OpenStudy (anonymous):

What is the simplified form of the expression 4(sqrt)v=4(sqrt)v6-(sqrt)25v-(sqrt)9v6

OpenStudy (anonymous):

\[4\sqrt{v}=4\sqrt{v^6}-\sqrt{25v}-\sqrt{9v^6}\]\[4\sqrt{v}=4\sqrt{v^6}-5\sqrt{v}-3\sqrt{v^6}\]\[4\sqrt{v}+5\sqrt{v}=\sqrt{v^6}\]\[4\sqrt{v}+5\sqrt{v}=\sqrt{v^6}\]\[9\sqrt{v}=v^3\]\[\sqrt{v}=\frac{1}{9}v^3\]\[v=\frac{1}{81}v^6\]\[81=v^5\]\[81=v^5\]\[v=81^\frac{1}{5}\]

OpenStudy (anonymous):

thank you